How would I solve this word problem?
Katy was given $35,000 when she retired. She invested some at 7% interest and the rest at 9% interest. If she earned $2,910 in interest in one year, how much was invested in each account?

Answer by ewatrrr(24785)

 
Hi,
Let x represent amount at 9%
.09x + .07(35000-x) = $2910
x = (2910 - 2450)/.02
x = 460/.02 = $23,000 at 9% and $12,000 at 7% 35000-23000+=+12000
CHECKING our answer***
$2070 + $840 = $2910
Wish You the Best in your Studies.